You’re taking your kid sister trick or treating.If you know that an average house gives out 6 pieces of candy with a standard deviation 2.3. What is the probability that your kid sister will hit up a house giving out 9 pieces pf candy?
1
Expert's answer
2019-11-12T12:35:19-0500

Let "X=" the number of pieces of candy: "X\sim B(n,p)."

Then


"mean=\mu_x=np""\sigma_x^2 =np(1-p)"

"{\sigma_x^2 \over \mu_x}={np(1-p) \over np}=1-p"

Given that "\mu_x=6, \sigma=2.3"


"1-p={(2.3)^2 \over 6}={5.29 \over 6}""p=1-{5.29 \over 6}={0.71 \over 6}"

"n={\mu_x \over p}\approx51"

"P(X=x)=\binom{n}{x}p^x(1-p)^{n-x}"

"P(X=9)\approx\binom{51}{9}({0.71 \over 6})^9({5.29 \over 6})^{51-9}\approx0.0698"

The probability that your kid sister will hit up a house giving out 9 pieces of candy is "0.0698"